# Kolmogorov-size particles in homogeneous and isotropic turbulence

### v1

@@ 1.1 The Maxey-Riley-Gatignol (MRG) Equation
In particle-laden turbulent flows the turbulent scales of the fluid phase are coupled in a non trivial manner with the solid phase. Properly resolving the flow around each particle is thus crucial to capture the fluid-solid interaction and describe the dynamics of the particles. Due to the prohibitive computational cost, however, most of the theoretical and numerical studies rely on approximations and models. As an example, in the context of particle clustering and preferential sampling — i.e. the tendency of particles to explore flow regions with specific properties — we mention the recent theoretical works by Goto & Vassilicos (2008), Coleman & Vassilicos (2009), Bragg et al. (2015) and Matsuda et al. (2024). Most of the models used are based on the seminal works of Maxey & Riley (1983) and Gatignol (1983), where the equation for small rigid spheres in a non-uniform flow (hereafter referred to as MRG equation) has been derived by exploiting linear perturbation theory. In these models, each particle is treated as a mathematical point source of mass, momentum and energy. In point-particle models, particles are indeed assumed to be much smaller than any structure of the flow, as the MRG equation holds when the fluid velocity field does not show a turbulent behaviour at the particle scale. In other words, the Reynolds number based on the particle diameter and the particle-fluid relative velocity has to be small, i.e. $\mathcal{O}(1)$.
@@
The models based on the MRG equation do not resolve the flow around the particles, and the influence of the solid phase on the fluid phase has to be modelled; see for example Ferrante & Elghobashi (2003), Gualtieri et al. (2015) and Vreman (2016). However, the back-reaction of the particles on the carrier flow and the inter-particle collisions are usually negligible in the limit of very dilute regimes with $\Phi_{V}=V_{s}/(V_{s}+V_{f})\leq 10^{-5}$ (where $\Phi_{V}$ is the volume fraction, and $V_{s}$ and $V_{f}$ are the volumes of the solid and fluid phases, respectively), small particles $D_{p}\leq\eta$ (where $\eta$ is the turbulent Kolmogorov scale) and small solid-to-fluid density ratios $\rho_{p}/\rho_{f}$ (Brandt & Coletti, 2022). In this case, the influence of the solid phase on the carrier fluid can be neglected and the fluid-particle interaction is often modelled with one-way coupling models, in which the particles move under the action of the flow, but they do not modulate it. Starting from the MRG equation, several corrections have been proposed over the years to account for several effects, and extend its range of validity to a wider range of parameters. For example, Saffman (1965) introduced a lift force which is crucial to properly model the particle dynamics in the presence of a linear shear. This force has been later extended by Mei (1992) to fit the numerical data of McLaughlin (1991) at large Reynolds numbers. Other corrections have been introduced to account for finite values of the particle Reynolds number. For example, we refer to the corrections to the drag term reported in Balachandar (2009), and to the different convolutional kernels for the Basset time-history force contribution proposed by Mei & Adrian (1992).
@@
The point particle approximation coupled with Direct Numerical Simulations of the Navier–Stokes equations (PP-DNS) has been widely used to investigate turbulent particle-laden flow in the one-way coupling regime in several scenarios (see Balachandar & Eaton, 2010, and references therein). Despite the large number of studies based on PP-DNS, however, clear understanding of the range of validity of the underlying assumptions of the point-particle model is not yet present, and studies that investigate its limitations are needed. In this respect, a step forward has been done in the last years thanks to the introduction of several numerical methods (PR-DNS) which resolve the flow around each particle by coupling the direct numerical simulations of the Navier–Stokes equations with the immersed boundary method (IBM); see for example Kajishima et al. (2001), Uhlmann (2005), Huang et al. (2007), Breugem (2012), Kempe & Fröhlich (2012) and Hori et al. (2022). Unlike experiments, PR-DNS allows to explain the underlying physical mechanisms and investigate the limitations of the point-particle approximation.
@@
In the context of forced homogeneous isotropic turbulence (HIT) with fixed particles, Burton & Eaton (2005) and Vreman (2016) found that, when compared with PR-DNS, the two-way coupled PP-DNS based on the Schiller-Naumann drag correlation captures fairly well the turbulence attenuation and the fraction of the turbulence dissipation rate due to the particles. Surprisingly, they found a good agreement also for $D_{p}/\eta=1.5$ (Burton & Eaton, 2005) and $D_{p}/\eta\approx 2.2$ (Vreman, 2016), although the model is expected to largely under-perform for these particle sizes. Mehrabadi et al. (2018) used PR-DNS to assess the validity of PP-DNS in a decaying isotropic turbulent particle-laden flow, focusing on the particle acceleration model. They found that the predictions of the PP-DNS models they considered are in excellent agreement compared with PR-DNS for small Stokes numbers. For large Stokes numbers, however, they found that PP-DNS under-predicts the true particle acceleration and that second moment quantities are not properly captured. They showed that the predictions improve once considering finite Reynolds number corrections to the model. Costa et al. (2020) tested the one-way point-particle approximation in a turbulence channel flow laden with small inertial particles, with high particle-to-fluid density ratios. They considered a volume fraction of $\Phi_{V}\approx 10^{-5}$ to ensure that the feedback of the particles on the fluid phase was negligible. They found that in the bulk of the channel the model predicts fairly well the statistic of the particles velocity. Close to the wall, however, they observed that the model fails, as it is not able to capture the shear-induced lift force acting on the particles, which instead is well predicted by the lift force model introduced by Saffman (1965).
@@
In this work, we further address the limit of the one-way point-particle approximation, and we use PR-DNS to investigate the reliability of one-way coupled PP-DNS in the context of forced homogeneous isotropic turbulence laden with Kolmogorov-size particles.
@@ 1.2 Particles in homogeneous isotropic turbulence
The majority of the numerical studies dealing with small particles $D_{p}\ll\eta$ are based on the point-particle approximation, as particle-resolved methods are prohibitively expensive in this case. Although the related model error is not known, the available numerical studies have shown that small particles may either amplify or damp the turbulence of the carrier flow, in agreement with previous experimental studies (Gore & Crowe, 1989). Squires & Eaton (1990) used two-way coupled PP-DNS to investigate homogeneous isotropic turbulence laden with small heavy spherical particles at $Re_{\lambda}=u^{\prime}\lambda/\nu\approx 38$ (where $u^{\prime}$ is the average velocity fluctuation, $\lambda$ is the Taylor length scale, and $\nu$ is the kinematic fluid viscosity). Compared to the particle-free case, they reported a significant attenuation of the fluid kinetic energy and of the dissipation rate. By looking at the energy spectrum, they found that the addition of the particles results into a relative enhancement of the energy at the small scales compared to the energy content at the large scales. They also showed that heavier particles cause a less selective modification of the turbulence properties. Heavy particles are indeed more uniformly dispersed by the turbulence, and cause a more homogeneous modification of the flow properties compared to lighter particles, that instead show a stronger preferential collection in regions of low vorticity and high strain (Maxey, 1987). Elghobashi & Truesdell (1993) used two-way coupled PP-DNS to investigate decaying homogeneous isotropic turbulence laden with small spherical particles. Besides confirming the non-uniform modulation of the energy spectrum, they observed that the energy enhancement at the small scales is accompanied by an increase of the viscous dissipation rate and, thus, by an enhancement of the rate of energy transfer from larger to smaller scales. Boivin et al. (1998) made use of two-way coupled PP-DNS to study the influence of small and heavy particles on forced homogeneous isotropic turbulence at $Re_{\lambda}=62$. They reported that the influence of the particles changes with their inertia, with the small-scale energy content being attenuated/enhanced by large/small particles. By investigating the spectrum of the fluid-particle energy exchange rate, they observed that particles act as a sink of kinetic energy at large scales, while they add kinetic energy to turbulence at the smallest scales. The large scale motions of the fluid drag the particles, while the small-scale fluctuations are driven by the presence of the solid phase. Druzhinin (2001) investigated the influence of small and heavy particles in decaying isotropic homogeneous turbulence at the initial Reynolds number of $Re_{\lambda}=30$ and $Re_{\lambda}=50$, with a focus on particles with very small inertia and small relaxation time. Unlike the previous works, they found that the turbulent kinetic energy and the viscous dissipation rate increase at all times compared to the particle-free case. The presence of the particles, indeed, largely enhances the small-scale energy content, while slightly reduces the large-scale energy content, with positive integral variation. This was later confirmed by Ferrante & Elghobashi (2003), that investigated by PP-DNS the influence of particles on decaying homogeneous isotropic turbulence with an initial Reynolds number of $Re_{\lambda}=75$.
@@
For large particles with $D_{p}>\eta$ the numerical studies are based on PR-DNS, as in this case the point-particle approximation does not hold. Lucci et al. (2010) and Lucci et al. (2011) investigated the influence of particles with size $D_{p}\approx\lambda$ (where $\lambda$ is the Taylor length scale) in decaying homogeneous turbulence. They observed that in contrast to what happens when $D_{p}<\eta$, the presence of the particles damps the turbulent kinetic energy of the fluid compared to the particle-free case at all times, and that the two-way coupling rate of change is always positive. Ten Cate et al. (2004) investigated the influence of particles with a solid-to-fluid density ratio of $\rho_{p}/\rho_{f}=1.15$ and $1.73$ on forced homogeneous isotropic turbulence at $Re_{\lambda}=u^{\prime}\lambda/\nu=61$, varying the volume fraction between $\Phi_{V}=0.02$ and $\Phi_{V}=0.1$. They found that the energy spectrum is enhanced for wavenumbers $\kappa>\kappa_{p}\approx 0.75\kappa_{D}$, where $\kappa_{D}=2\pi/D_{p}$, while it is attenuated for $\kappa<\kappa_{p}$. These results were later confirmed by Yeo et al. (2010) at $Re_{\lambda}\approx 60$, $\rho_{p}/\rho_{f}=1.4$ and $\Phi_{V}=0.06$. Uhlmann & Chouippe (2017) considered particles with $D_{p}/\eta\approx 5-8$ and $\rho_{p}/\rho_{f}=1.5$ at the larger Reynolds number of $Re_{\lambda}\approx 130$. They focused on the dynamics of the particles, and observed that finite-size inertial particles exhibit a moderate level of clustering, as later confirmed also by Chiarini & Rosti (2024). Olivieri et al. (2022a) considered the effect of particles on homogeneous isotropic turbulence at the larger Reynolds number of $Re_{\lambda}\approx 400$, which ensures a well developed inertial range of scales. They set the volume fraction at $\Phi_{V}=0.079$, and investigated the turbulence modulation by particles with size $D_{p}/\eta=123$ and solid-to-fluid density ratio between $1.3\leq\rho_{p}/\rho_{f}\leq 100$. They showed that the solid phase modifies the energy cascade described by Richardson and Kolmogorov; the fluid-solid coupling drives the energy cascade at large scales, while the classical energy cascade is restored at scales smaller than the particle size. Oka & Goto (2022) studied the turbulence modulation due to spherical particles with $7.8\leq D_{p}/\eta\leq 64$ by setting the volume fraction at $\Phi_{V}=8.1\times 10^{-3}$ and the reference Reynolds number at $Re_{\lambda}\leq 100$. They found that the turbulent kinetic energy content monotonically decreases with $D_{p}$, due to the increase of the energy dissipation rate in the wake of the particles. More recently, Chiarini et al. (2024) and Chiarini & Rosti (2024) investigated by PR-DNS how the flow modulation changes with $D_{p}$ and $\rho_{p}$. They set the Reynolds number to $Re_{\lambda}\approx 400$ and the volume fraction to $\Phi_{V}=0.079$, and varied the particles size and the solid-to-fluid density ratio in the $16\leq D_{p}/\eta\leq 123$ and $1.3\leq\rho_{p}/\rho_{f}\leq 100$ range. Chiarini et al. (2024) found that, in presence of an inhomogeneous mean shear, particles might modulate the largest scales of the flow towards an anisotropic state, while preserving homogeneity and isotropy at the smaller scales. Chiarini & Rosti (2024) observed that interface-resolved particles enhance flow intermittency favouring events with large localised velocity gradients. For the smallest and heaviest particles, they found that the classical energy cascade is subdominant at all scales, and that the energy transfer is completely driven by the fluid-solid coupling term. Cannon et al. (2024) investigated the effect of the Reynolds number on the flow modulation by finte-size particles in homogeneous isotropic turbulence. Notably, they observed that the modulation of the turbulent kinetic energy has little dependence on $Re_{\lambda}$, and that particles modulate turbulence also at the smallest Reynolds numbers.
@@
While a relatively larger body of literature has investigated the dynamics of particle of size larger and smaller than Kolmogorov size, fewer works have considered Kolmogorov-size particles with $D_{p}\approx\eta$, which are the focus of the present work. From an experimental point of view the $D_{p}\approx\eta$ case is complex, as it requires a resolution of sub-Kolmogorov scales when measuring the velocity perturbations near the particles (Tanaka & Eaton, 2010). Numerical schemes based on the MRG equation, which are commonly used for $D_{p}\ll\eta$, are generally thought of as not valid when $D_{p}\approx\eta$ (Balachandar & Eaton, 2010). PR-DNS, on the other hand, becomes prohibitively expensive as $D_{p}$ decreases when $Re$ is sufficiently large, due to the extra resolution required to properly resolve both the flow perturbations induced by the particles and all the turbulence scales. Among the few works available, we mention Hwang & Eaton (2006) that experimentally investigated the influence of a dilute dispersion of particles with $D_{p}\approx\eta$ in forced homogeneous isotropic turbulence at $Re_{\lambda}\approx 230$. They observed that Kolmogorov-size particles attenuate the turbulent global kinetic energy and the viscous dissipation rate up to $40\%$ and $50\%$ for a mass loading $\Phi_{M}=\rho_{p}V_{p}/(\rho_{f}V_{f})$ of $\Phi_{M}=0.3$. Schneiders et al. (2017) studied by PR-DNS the interaction of decaying isotropic turbulence with finite-size $D_{p}\approx\eta$ particles. They set the initial Reynolds number of the flow at $Re_{\lambda}=79$, and varied the solid-to-fluid density ratio between $40\leq\rho_{p}/\rho_{f}\leq 5000$ and the mass loading between $0.01\leq\Phi_{M}\leq 1$. They observed that in the vicinity of the particles the viscous dissipation rate of the fluid is amplified due to the large velocity gradients that are generated by the boundary conditions at the surface of the particles (see also Tanaka & Eaton, 2010; Brändle de Motta et al., 2016; Chiarini & Rosti, 2024). Particles also release kinetic energy to the fluid by locally accelerating the surrounding flow, similarly to what seen by Chiarini & Rosti (2024) for larger particles. From a global viewpoint, Schneiders et al. (2017) observed that, for large $\rho_{p}/\rho_{f}$, particles with $D/\eta\approx 1$ induce local velocity disturbances that significantly modulate the distribution and the decay of the fluid kinetic energy at all scales.

### v2

@@ 1.1 The Maxey-Riley-Gatignol (MRG) Equation
In particle-laden turbulent flows the turbulent scales of the fluid phase are coupled in a non trivial manner with the solid phase. Properly resolving the flow around each particle is thus crucial to capture the fluid-solid interaction and describe the dynamics of the particles. Due to the prohibitive computational cost, however, most of the theoretical and numerical studies rely on approximations and models. As an example, in the context of particle clustering and preferential sampling — i.e. the tendency of particles to explore flow regions with specific properties — we mention the recent theoretical works by Goto & Vassilicos (2008), Coleman & Vassilicos (2009), Bragg et al. (2015) and Matsuda et al. (2024). Most of the models used are based on the seminal works of Maxey & Riley (1983) and Gatignol (1983), where the equation for small rigid spheres in a non-uniform flow (hereafter referred to as MRG equation) has been derived by exploiting linear perturbation theory. In these models, each particle is treated as a mathematical point source of mass, momentum and energy. In point-particle models, particles are indeed assumed to be much smaller than any structure of the flow, as the MRG equation holds when the fluid velocity field does not show a turbulent behaviour at the particle scale. In other words, the Reynolds number $Re_{p}$ based on the particle diameter and the particle-fluid relative velocity has to be small, i.e. $Re_{p}\ll 1$.
@@
The models based on the MRG equation do not resolve the flow around the particles, and the influence of the solid phase on the fluid phase has to be modelled; see for example Ferrante & Elghobashi (2003), Gualtieri et al. (2015) and Vreman (2016b). However, the back-reaction of the particles on the carrier flow and the inter-particle collisions are usually negligible in the limit of very dilute regimes with $\Phi_{V}=V_{s}/(V_{s}+V_{f})\leq 10^{-5}$ (where $\Phi_{V}$ is the volume fraction, and $V_{s}$ and $V_{f}$ are the volumes of the solid and fluid phases, respectively), small particles $D_{p}<\eta$ (where $\eta$ is the turbulent Kolmogorov scale) and small solid-to-fluid density ratios $\rho_{p}/\rho_{f}$ (Brandt & Coletti, 2022). In this case, the influence of the solid phase on the carrier fluid can be neglected and the fluid-particle interaction is often modelled with one-way coupling models, in which the particles move under the action of the flow, but they do not modulate it. Starting from the MRG equation, several corrections have been proposed over the years to account for several effects, and extend its range of validity to a wider range of parameters. For example, Saffman (1965) introduced a lift force which is crucial to properly model the particle dynamics in the presence of a linear shear. This force has been later extended by Mei (1992) to fit the numerical data of McLaughlin (1991) at large Reynolds numbers. Other corrections have been introduced to account for finite values of the particle Reynolds number. For example, we refer to the corrections to the drag term reported in Balachandar (2009), and to the different convolutional kernels for the Basset time-history force contribution proposed by Mei & Adrian (1992).
@@
The point particle approximation coupled with Direct Numerical Simulations of the Navier–Stokes equations (PP-DNS) has been widely used to investigate turbulent particle-laden flow in the one-way coupling regime in several scenarios (see Balachandar & Eaton, 2010, and references therein). Despite the large number of studies based on PP-DNS, however, clear understanding of the range of validity of the underlying assumptions of the point-particle model is not yet present, and studies that investigate its limitations are needed. In this respect, a step forward has been done in the last years thanks to the introduction of several numerical methods (PR-DNS) which couple the Direct Numerical Simulations of the Navier–Stokes equations with techniques that resolve the flow around each particle and capture the effect of the no-slip boundary condition at the particles’ surface on the flow. We mention the arbitrary Lagrangian-Eulerian (ALE) technique (Hu et al., 2001), the overset-grid approaches (Burton & Eaton, 2005; Koblitz et al., 2017; Vreman, 2016a; Horne & Mahesh, 2019), the Physalis technique introduced by Prosperetti & Oguz (2001) and the immersed boundary method (IBM) (Kajishima et al., 2001; Uhlmann, 2005; Huang et al., 2007; Breugem, 2012; Kempe & Fröhlich, 2012; Hori et al., 2022). Unlike experiments, PR-DNS allows to explain the underlying physical mechanisms and investigate the limitations of the point-particle approximation.
@@
In the context of forced homogeneous isotropic turbulence (HIT) with fixed particles, Burton & Eaton (2005) and Vreman (2016b) found that, when compared with PR-DNS, the two-way coupled PP-DNS based on the Schiller-Naumann drag correlation captures fairly well the turbulence attenuation and the fraction of the turbulence dissipation rate due to the particles. Surprisingly, they found a good agreement also for $D_{p}/\eta=1.5$ (Burton & Eaton, 2005) and $D_{p}/\eta\approx 2.2$ (Vreman, 2016b), although the model is expected to largely under-perform for these particle sizes. Mehrabadi et al. (2018) used PR-DNS to assess the validity of PP-DNS in a decaying isotropic turbulent particle-laden flow, focusing on the particle acceleration model. They found that the predictions of the PP-DNS models they considered are in excellent agreement compared with PR-DNS for small Stokes numbers. For large Stokes numbers, however, they found that PP-DNS under-predicts the true particle acceleration and that second moment quantities are not properly captured. They showed that the predictions improve once considering finite Reynolds number corrections to the model. Costa et al. (2020) tested the one-way point-particle approximation in a turbulent channel flow laden with small inertial particles, with high particle-to-fluid density ratios. They considered a volume fraction of $\Phi_{V}\approx 10^{-5}$ to ensure that the feedback of the particles on the fluid phase was negligible. They found that in the bulk of the channel the model predicts fairly well the statistics of the particles velocity. Close to the wall, however, they observed that the model fails, as it is not able to capture the shear-induced lift force acting on the particles, which instead is well predicted by the lift force model introduced by Saffman (1965).
@@
In this work, we further address the limit of the one-way point-particle approximation, and we use PR-DNS to investigate the reliability of one-way-coupled PP-DNS in the context of forced homogeneous isotropic turbulence laden with Kolmogorov-size particles.
@@ 1.2 Particles in homogeneous isotropic turbulence
The majority of the numerical studies dealing with small particles $D_{p}\ll\eta$ are based on the point-particle approximation, as particle-resolved methods are prohibitively expensive in this case. Although the related model error is not known, the available numerical studies have shown that small particles may either amplify or damp the turbulence of the carrier flow, in agreement with previous experimental studies (Gore & Crowe, 1989). Squires & Eaton (1990) used two-way-coupled PP-DNS to investigate homogeneous isotropic turbulence laden with small heavy spherical particles at $Re_{\lambda}=u^{\prime}\lambda/\nu\approx 38$ (where $u^{\prime}$ is the average velocity fluctuation, $\lambda$ is the Taylor length scale, and $\nu$ is the kinematic fluid viscosity). Compared to the particle-free case, they reported a significant attenuation of the fluid kinetic energy and of the dissipation rate. By looking at the energy spectrum, they found that the addition of the particles results in a relative enhancement of the energy at the small scales compared to the energy content at the large scales. They also showed that heavier particles cause a less selective modification of the turbulence properties. Heavy particles are indeed more uniformly dispersed by the turbulence, and cause a more homogeneous modification of the flow properties compared to lighter particles, that instead show a stronger preferential collection in regions of low vorticity and high strain (Maxey, 1987). Elghobashi & Truesdell (1993) used two-way coupled PP-DNS to investigate decaying homogeneous isotropic turbulence laden with small spherical particles. Besides confirming the non-uniform modulation of the energy spectrum, they observed that the energy enhancement at the small scales is accompanied by an increase of the viscous dissipation rate and, thus, by an enhancement of the rate of energy transfer from larger to smaller scales. Boivin et al. (1998) made use of two-way-coupled PP-DNS to study the influence of small and heavy particles on forced homogeneous isotropic turbulence at $Re_{\lambda}=62$. They reported that the influence of the particles changes with their inertia, with the small-scale energy content being attenuated/enhanced by large/small particles. By investigating the spectrum of the fluid-particle energy exchange rate, they observed that particles act as a sink of kinetic energy at large scales, while they add kinetic energy to turbulence at the smallest scales. The large scale motions of the fluid drag the particles, while the small-scale fluctuations are driven by the presence of the solid phase. Druzhinin (2001) investigated the influence of small and heavy particles in decaying isotropic homogeneous turbulence at the initial Reynolds number of $Re_{\lambda}=30$ and $Re_{\lambda}=50$, with a focus on particles with very small inertia and small relaxation time. Unlike the previous works, they found that the turbulent kinetic energy and the viscous dissipation rate increase at all times compared to the particle-free case. The presence of the particles, indeed, largely enhances the small-scale energy content, while slightly reduces the large-scale energy content, with positive integral variation. This was later confirmed by Ferrante & Elghobashi (2003), that investigated by PP-DNS the influence of particles on decaying homogeneous isotropic turbulence with an initial Reynolds number of $Re_{\lambda}=75$.
@@
For large particles with $D_{p}>\eta$ the numerical studies are based on PR-DNS, as in this case the point-particle approximation does not hold. Lucci et al. (2010) and Lucci et al. (2011) investigated the influence of particles with size $D_{p}\approx\lambda$ in decaying homogeneous turbulence. They observed that in contrast to what happens when $D_{p}<\eta$, the presence of the particles damps the turbulent kinetic energy of the fluid compared to the particle-free case at all times, and that the two-way coupling rate of change is always positive. Ten Cate et al. (2004) investigated the influence of particles with a solid-to-fluid density ratio of $\rho_{p}/\rho_{f}=1.15$ and $1.73$ on forced homogeneous isotropic turbulence at $Re_{\lambda}=u^{\prime}\lambda/\nu=61$, varying the volume fraction between $\Phi_{V}=0.02$ and $\Phi_{V}=0.1$. They found that the energy spectrum is enhanced for wavenumbers $\kappa>\kappa_{p}\approx 0.75\kappa_{D}$, where $\kappa_{D}=2\pi/D_{p}$, while it is attenuated for $\kappa<\kappa_{p}$. These results were later confirmed by Yeo et al. (2010) at $Re_{\lambda}\approx 60$, $\rho_{p}/\rho_{f}=1.4$ and $\Phi_{V}=0.06$. Uhlmann & Chouippe (2017) considered particles with $D_{p}/\eta\approx 5-8$ and $\rho_{p}/\rho_{f}=1.5$ at the larger Reynolds number of $Re_{\lambda}\approx 130$. They focused on the dynamics of the particles, and observed that finite-size inertial particles exhibit a moderate level of clustering, as later confirmed also by Chiarini & Rosti (2024). Olivieri et al. (2022a) considered the effect of particles on homogeneous isotropic turbulence at the larger Reynolds number of $Re_{\lambda}\approx 400$, which ensures a well developed inertial range of scales. They set the volume fraction at $\Phi_{V}=0.079$, and investigated the turbulence modulation by particles with size $D_{p}/\eta=123$ and solid-to-fluid density ratio between $1.3\leq\rho_{p}/\rho_{f}\leq 100$. They showed that the solid phase modifies the energy cascade described by Richardson and Kolmogorov; the fluid-solid coupling drives the energy cascade at large scales, while the classical energy cascade is restored at scales smaller than the particle size. Oka & Goto (2022) studied the turbulence modulation due to spherical particles with $7.8\leq D_{p}/\eta\leq 64$ by setting the volume fraction at $\Phi_{V}=8.1\times 10^{-3}$ and the reference Reynolds number at $Re_{\lambda}\leq 100$. They found that the turbulent kinetic energy content monotonically decreases with $D_{p}$, due to the increase of the energy dissipation rate in the wake of the particles. More recently, Chiarini et al. (2024) and Chiarini & Rosti (2024) investigated by PR-DNS how the flow modulation changes with $D_{p}$ and $\rho_{p}$. They set the Reynolds number to $Re_{\lambda}\approx 400$ and the volume fraction to $\Phi_{V}=0.079$, and varied the particles size and the solid-to-fluid density ratio in the $16\leq D_{p}/\eta\leq 123$ and $1.3\leq\rho_{p}/\rho_{f}\leq 100$ range. Chiarini & Rosti (2024) observed that interface-resolved particles enhance flow intermittency favouring events with large localised velocity gradients. For the smallest and heaviest particles, they found that the classical energy cascade is subdominant at all scales, and that the energy transfer is completely driven by the fluid-solid coupling term. Cannon et al. (2024) investigated the effect of the Reynolds number on the flow modulation by finte-size particles in homogeneous isotropic turbulence. Notably, they observed that the modulation of the turbulent kinetic energy has little dependence on $Re_{\lambda}$, and that particles modulate turbulence also at the smallest Reynolds numbers.
@@
While a relatively larger body of literature has investigated the dynamics of particle of size larger and smaller than Kolmogorov size, few works have considered Kolmogorov-size particles with $D_{p}\approx\eta$, which are the focus of the present work. From an experimental point of view the $D_{p}\approx\eta$ case is complex, as it requires a resolution of sub-Kolmogorov scales when measuring the velocity perturbations near the particles (Tanaka & Eaton, 2010). Numerical schemes based on the MRG equation, which are commonly used for $D_{p}\ll\eta$, are generally thought of as not valid when $D_{p}\approx\eta$ (Balachandar & Eaton, 2010). PR-DNS, on the other hand, becomes prohibitively expensive as $D_{p}$ decreases when $Re$ is sufficiently large, due to the extra resolution required to properly resolve both the flow perturbations induced by the particles and all the turbulence scales. Among the few works available, we mention Hwang & Eaton (2006) that experimentally investigated the influence of a dilute dispersion of particles with $D_{p}\approx\eta$ in forced homogeneous isotropic turbulence at $Re_{\lambda}\approx 230$. They observed that Kolmogorov-size particles attenuate the turbulent global kinetic energy and the viscous dissipation rate up to $40\%$ and $50\%$ for a mass loading $\Phi_{M}=\rho_{p}V_{p}/(\rho_{f}V_{f})$ of $\Phi_{M}=0.3$. Yang & Shy (2005) experimentally investigated the settling of small solid particles in homogeneous isotropic turbulence, with $\rho_{p}\gg\rho_{f}$ and $D_{p}\leq\eta$ such that $Re_{p}<1$ and $0\leq St\leq 2$. They observed that at these parameters the tendency of particles to form clusters is maximum for $St\approx 1$, and that the clusters are distributed along the periphery of intense vortical structures. Poelma et al. (2007), instead, investigated the influence of $D/\eta=O(1)$ and $\rho_{p}/\rho_{f}=O(1)$ particles on the decay rate of grid generated turbulence. According to their experiments, the presence of the particles moves the onset of the turbulence decay upstream, and promotes the flow anisotropy. For recent experimental works concerning sub-Kolmogorov $D_{p}<\eta$ particles, we refer the reader to Sumbekova et al. (2017); Petersen et al. (2019); Hassaini et al. (2023) (clustering) and to Hassaini & Coletti (2022) (turbulence modification). Schneiders et al. (2017) studied by PR-DNS the interaction of decaying isotropic turbulence with finite-size $D_{p}\approx\eta$ particles. They varied the solid-to-fluid density ratio between $40\leq\rho_{p}/\rho_{f}\leq 5000$ and the mass loading between $0.01\leq\Phi_{M}\leq 1$. They set the Reynolds number of the flow at $Re_{\lambda}(t_{0})=79$, a starting value for which the flow lacks a well-defined inertial range. They observed that in the vicinity of the particles the viscous dissipation rate of the fluid is amplified due to the large velocity gradients that are generated by the boundary conditions at the surface of the particles (see also Tanaka & Eaton, 2010; Brändle de Motta et al., 2016; Chiarini & Rosti, 2024). Particles also release kinetic energy to the fluid by locally accelerating the surrounding flow, similarly to what seen by Chiarini & Rosti (2024) for larger particles. From a global viewpoint, Schneiders et al. (2017) observed that, for large $\rho_{p}/\rho_{f}$, particles with $D/\eta\approx 1$ induce local velocity disturbances that significantly modulate the distribution and the decay of the fluid kinetic energy at all scales. Overall, despite the interest an exhaustive characterisation of how Kolmogorov-size particles modulate turbulence at a Reynolds number that is large enough to ensure a proper separation of scales is still lacking.